NERSCPowering Scientific Discovery Since 1974

June 7

Attending:  Eric, Jay and Katie from NERSC and users Andrei, Jeff A. and Jeff P.

Utilization:  Cluster tends to be running at about 50% of capacity on average the past month.

Recent Downtimes/Outages:  6/1:  SWO for most of the day:  /home migration to new hardware was completed, updated BestMan on pdsfdtn1.  This upgrade will hopefully address some timeout bug that had affected STAR transfers.

Upcoming Downtimes:  No clusterwide outages planned but eliza3 needs a reboot.  This shouldn't require a downtime but might degrade performance for a while.

Question:  Are we on new /common yet?  Answer:  No, has to do with investigations of using moab instead of maui with torque.

Procurements:  Dell has been difficult with the Alice procurements because they want some changes which are difficult for us to make so it's been delayed a bit.

Project Accounts:  Eric reported that a working project account is in place an is being test.  Will also need to test logging w/security to be sure we record what we need.

Torque:  Eric reports that he's working on webpages and a pbsusers script, similar to the old sgeusers script is available.

Grid Transfers:  Andrei reports that ATLAS transfers are still working, at least for small files, but there is a problem with deleting files that therer is a ticket about.  He say this isn't realy high priority but they will want to get it working.

Andrei reports reading on a new cvmfs mailing list that a cache size of about 50GB works well and that is what we had decided months ago.